OpenStudy (anonymous):

What is the completion to sin^2theta+tan^2theta+cos^2theta?

OpenStudy (anonymous):

it's all about knowing your identities! \[\sin^2(\theta)+\cos^2(\theta)=1\] is one identity use that and you're left with \[1+\tan^2(\theta)= ?\]

OpenStudy (anonymous):

oh look, another trig identity! \[\tan^2(\theta)+1=\sec^2(\theta)\] and then you're done

OpenStudy (anonymous):

Ohh ok. Thank you!